OpenStudy (anonymous):

At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

if i remember correctly, if there are \(n\) people there are \[\binom{n}{2}\] handshakes

OpenStudy (misty1212):

set \[\frac{n(n-1)}{2}=66\] and solve for \(n\)

OpenStudy (anonymous):

This is a bit confusing...

OpenStudy (misty1212):

i can see that

OpenStudy (misty1212):

if \[\frac{n(n-1)}{2}=66\] then \[n(n-1)=132\]

OpenStudy (misty1212):

see if you can think of two consecutive numbers whose product is 132

OpenStudy (misty1212):

you can probably guess it quickly

OpenStudy (ybarrap):

 Notice the sequence: 1,3,6,10,15 Let's see how they are related:  Note that the difference between the number of handshakes for each step is equal to the number of people doing the handshaking: $$ a_1-a_0=1-0=1\\ a_2-a_1=3-1=2\\ a_3-a_2=6-3=3\\ a_4-a_3=10-6=4\\ a_5-a_4=15-10=5\\ \cdots\\ a_n-a_{n-1}=n $$ If you add everything on the left side of the equal sign and everything on the right side of the equal sign, you get: $$ a_n-a_0=\cfrac{n(n+1)}{2}\\ a_n=\cfrac{n(n+1)}{2}\\ $$ We want \(a_n=66\): $$ 66=\cfrac{n(n+1)}{2}\\\\ 2\times 66=n^2+n\\ n^2+n-132=0\\ \implies n = 11 $$ So there will be 66 handshakes if there are 11 people at the party.